# FreeA: Human-object Interaction Detection using Free Annotation Labels

### v1

@@ Abstract
Recent human-object interaction (HOI) detection approaches rely on high cost of manpower and require comprehensive annotated image datasets. In this paper, we propose a novel self-adaption language-driven HOI detection method, termed as FreeA, without labeling by leveraging the adaptability of CLIP to generate latent HOI labels. To be specific, FreeA matches image features of human-object pairs with HOI text templates, and a priori knowledge-based mask method is developed to suppress improbable interactions. In addition, FreeA utilizes the proposed interaction correlation matching method to enhance the likelihood of actions related to a specified action, further refine the generated HOI labels. Experiments on two benchmark datasets show that FreeA achieves state-of-the-art performance among weakly supervised HOI models. Our approach is +8.58 mean Average Precision (mAP) on HICO-DET and +1.23 mAP on V-COCO more accurate in localizing and classifying the interactive actions than the newest weakly model, and +1.68 mAP and +7.28 mAP than the latest weakly+ model, respectively. Code will be available at https://drliuqi.github.io/.
Index Terms:
Human-object interaction detection, weakly supervised, object detection.
publicationid: pubid: $0$–$00$ 2024 IEEE
@@
I Introduction
Whether one-stage [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] or two-stage [1, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] HOI detection models, predominantly relies on computationally heavy training and requires extensive-annotation datasets (Figure 1(a) and Figure 1(b)). Taking the HICO-Det dataset as an example, the weakly+ or weakly supervised approaches need to annotate 117,871 interaction labels from (117,871$\times$600) or (117,871$\times$23) potential combinations. It is quite resource-intensive. As shown in Figure 1(b), HOI models can be categorized into twofolds [23]: weakly+ using $\langle$“interaction”, “object”$\rangle$ label  [24, 25, 26], e.g., eat-banana, and weakly with only $\langle$“interaction”$\rangle$ labels, e.g., eat. Although weakly supervised HOI detection models typically employs image-level labels, with only interaction action categories, they still require abundant annotations for large-scale datasets to achieve satisfactory performance. Compared to weakly+ supervised methods, weakly supervised ones are more challenging since they require to explore objects within a broader interaction space  [23]. For instance, when considering the action “riding”, there could be a wide range of objects involved. Weakly+ supervised models with $\langle$“ride”, “motorcycle”$\rangle$ labels effectively focus on motorcycles, while it is not easy for weakly supervised models with $\langle$“ride”$\rangle$ to determine the accurate object because of multiple matching choices, such as riding bicycle. The proposed method, called as FreeA, belongs to the latter [^1] The reason why FreeA is considered as one weakly supervised HOI model is that the used labels in FreeA are automatically generated by CLIP.. In addition, another sub-direction of research in HOI detection is zero-shot HOI detection [15, 27, 28, 29, 30, 18, 31, 32, 33, 34, 9, 35, 36], where the model is trained using only a subset of annotated datasets but is tested on interaction categories not seen during training. It can be observed that zero-shot HOI detection focuses on discovering unseen samples. However, it still relies on training with manually annotated labels.
Motivated by the ability of text CLIP (contrastive language image pretraining) estimates best paired with a given image, FreeA is developed for achieving HOI detection task without the need for manual labels (Figure 1(c)). The FreeA mainly comprises three-folds: candidate image construction (CIC), human-object potential interaction mining (PIM), and human-object interaction inference (HII). In the phase of CIC, FreeA is plug-and-play applying existing object detection methods for all potential instances localization, and utilizes spatial denoising and pairing techniques to establish candidate interaction pairs within the image. The PIM module extensively leverages the adaptability of the text-image model, i.e., CLIP, in the target domain to align the high-dimensional image features with HOI interaction templates, which generates the similarity vectors of candidate interaction relationships. Then, the HII module combines the resulting similarity vectors with a priori HOI action masks to mitigate interference from irrelevant relationships, and augments the likelihood of specific actions through the proposed interaction correlation matching method. Moreover, an adaptive threshold in HII is used to dynamically generate HOI labels for training.
Our key contributions are summarized as threefolds:
@@ I Introduction
    We propose a novel HOI detection method, namely FreeA, automatically generates HOI labels.
@@
    HOI detection includes various interactions among multiple instances. Three key challenges are required to tackle when using CLIP to generate labels, namely, multiple actions selection, filtering out irrelevant actions, and refining CLIP’s coarse labels. To address that, three corresponding modules are presented that significantly improve the effectiveness of the interactions localization and classification.
@@
    A broad variety of experiments are conducted to demonstrate the remarkable results of the proposal on HOI detection task, and ours performs the best among all weakly+, weakly and several fully supervised HOI models.
II Related Work
@@
Supervised HOI Detection. Supervised HOI models train their networks with the help of manually annotation $\langle$“person”, “object”, “action”$\rangle$. Their networks are presented as two-stage or single-stage pipelines. The two-stage methods first use pre-trained object detection network [37, 38, 39, 1] to detect humans and objects, and then pair them one by one into the interactive discrimination network to achieve HOI detection [1, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. However, it is pretty inefficient for one-to-one pairing between humans and objects [2, 8]. To address that, single-stage HOI detection based on transformer is gradually developed via end-to-end solution  [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. HOITrans [8] and QPIC [40] applied extractors and encoders in DETR [41] to extract features for global feature encoding. It had been verified that text information enables to improve the HOI detection performance [42, 9]. Current HOI approaches, e.g., GEN-VLKT [9] and HOICLIP [35], took use of CLIP [43] to train the encoding network via image-text pairs.
@@
Figure 2: Method overview. Starting from an existing HOI model, we apply the candidate image construction to extract humans and objects by detection and segmentation, and establish one-to-one human-object pairing. The human-object potential interaction mining module gets initial HOI interaction labels from candidate image pairs, and uses CLIP for domain adaptation. The human-object interaction inference module further refines these interaction labels by using prior knowledge-based mask to eliminate implausible actions and using interaction correlation matching method to enhance relevant actions similarity. Finally, HOI labels are generated for model training through dynamic threshold selector and interaction action filter.
@@
Weakly Supervised HOI Detection. Weakly supervised HOI detection generally uses image-level interaction labels for training [44, 24, 23]. They can be divided into twofolds: “weakly+” with $\langle$“interaction”, “object”$\rangle$ annotations [44, 24, 26], and “weakly” with only $\langle$“interaction”$\rangle$ annotations [23]. MX-HOI [24] proposed a momentum independent learning framework using weakly+ supervised ($\langle$“interaction”, “object”$\rangle$). Align-Former [44] proposed an “align layer” to achieve pseudo alignment for training based on transformer architecture. Nevertheless, these methods suffer from noisy human-object association and ambiguous interaction types [26]. Therefore, PGBL [26] used CLIP as an interactive prompt network and HOI knowledge to enhance interaction judgment at the instance level. Weakly-HOI [23] applied a language model to query for unnecessary interaction pairing reduction with image-level $\langle$“interaction”$\rangle$ annotations. Both of them, however, are still dependent on pre-annotated datasets at the cost of manpower. Unlike the latest methods [23] and [26], we propose a method for HOI detection that does not require manual annotation.
@@
Zero-shot HOI Detection. The goal of zero-shot HOI detection is to train on a subset of labels and test using another set of unseen labels to detect interactions that were not encountered during training. Many methods [15, 27, 28, 29, 30, 18, 31, 32, 33, 34, 9, 35, 36] are investigated to handle zero-shot HOI detection. Shen et al. [32] pioneered the application of zero-shot learning methods to address the long-tail problem in HOI detection. They introduced a decomposition model for HOI detection, separately inferring verbs and objects, enabling the detection of new verb-object combinations during testing. S2S [36] embedded both semantic and spatial information into the visual stream, proposing a person-object interaction detection based on verb-object relation reasoning. GEN-VLKT [9] is a simple yet effective framework that utilizes CLIP for knowledge transfer [35, 26, 23], thereby discovering unknown samples. The introduction of zero-shot learning enhances the adaptability of these methods to real-world scenarios. However, these methods still require manually annotated complete HOI labels.
III Method
@@
As shown in Figure 2, we propose a novel plug-and-play HOI framework, namely, FreeA, to reduce the requirements of annotations. The proposed framework includes candidate image construction (Sec. III-B), human-object potential relationship mining (Sec. III-C), and human-object relationship inference (Sec. III-D). Details are introduced below.
III-A HOI Model
Motivated by the success of GEN-VLKT [9], we use it as the HOI model shown in Figure 2, and training is initialized with the parameters of COCO trained DETR [41]. It is worth noting that a better HOI model can promote the overall effect of FreeA, as verified in Sec IV-B.
III-B Candidate Image Construction
@@
To achieve automatic labels generation, we need to accurately localize humans and objects before, where Yolov8222https://github.com/ultralytics/ultralytics model is used for localization. Given an input image $I$, we obtain a collection of instance bounding boxes $\bm{B}=\{(\bm{b}_{i}|i=1,2,...,N)\}$, where $\bm{b}_{i}=(x_{i},y_{i},W_{i},H_{i},c_{i})$ from Yolov8, with $x_{i}$ and $y_{i}$ representing the center coordinates of the $i$th bounding box. $W_{i}$ and $H_{i}$ are the width and height of the bounding box, and $c_{i}$ denotes the category of the instance within the box. Here, $N$ represents the number of detected instances.
@@
Subsequently, we trim the images using the bounding boxes and then pair humans with objects to create candidate images. $N_{I}$ candidate images are obtained, where $N_{I}=N_{h}\times N_{o}$, and $N_{h}+N_{o}=N$. $N_{h}$ and $N_{o}$ represent the number of humans and objects, respectively. To eliminate interference from redundant instances and background information, we apply instance segmentation to assist the PIM module to focus on interest of interactions.
III-C Human-object Potential Interaction Mining
@@
As shown in Figure 3, the PIM module aims to mine potential relationships between humans and objects. The pre-trained CLIP is used to transfer knowledge from source domain to target domain, where features are extracted from both texts and images. Then by computing cross-modal similarity, text-image pairs are matched.
@@
CLIP Image Encoder. The image encoder $\mathbf{F}_{IE}$ of CLIP is employed to process the set of $N_{I}$ candidate images, with the result of image encoding $\bm{I}_{E}\in\mathbb{R}^{N_{I}\times C_{IE}}$. The $\bm{C}_{IE}$ denotes the dimension of the image encoder, and $\bm{I}_{B}$ represents the collection of candidate images. We have:
Figure 3: Human-object potential interaction mining module. The candidate image and HOI action text prompt templates are separately fed into the image encoder and text encoder of CLIP to obtain high-level feature encodings. The image encoding, denoted as $I_{E}$, is paired with the text encodings $T_{E1}$ and $T_{E2}$ to calculate textual similarities. The similarity calculations are specified by Eq. 3 and 4.
@@
CLIP Text Encoder. We start with a text template creation, denoted as $\bm{T}_{1}$, in the format “a photo of a person verb-ing an object”. For example, the triplet $\langle$ “human”, “ride”, “motorcycle”$\rangle$ is transformed into “a photo of a person riding a motorcycle”. After that, $\bm{T}_{1}$ is input into the text encoder $\mathbf{F}_{TE}$, leading to a text information matrix $\bm{T}_{E1}$ $\in\mathbb{R}^{N_{T}\times C_{TE}}$. The $N_{T}$ represents the number of texts, i.e., the number of HOI interaction action categories, and $C_{TE}$ is the dimensionality of the encoded text.
To emphasize the importance of verbs in HOI relationships, another type of text template $\bm{T}_{2}$ in the format “a photo of a person verb-ing”, has been constructed. $\bm{T}_{1}$ and $\bm{T}_{2}$ are distinct text templates for different HOI actions, with corresponding to information matrices $\bm{T}_{E1}$ and $\bm{T}_{E2}$, respectively. They are formulated as:
@@
Figure 4: Details of the interaction correlation matching method. After selecting the action with the highest similarity, we perform action correlation extraction. For example, if “race” is the selected action, related actions such as “ride”, “straddle”, and “sit” will be identified. We then amplify the similarities of these related actions, making actions highly correlated with “race” more likely to be chosen in the subsequent steps.
III-D Human-object Interaction Inference
@@
Interaction Correlation Matching (ICM). If a certain action occurs, other actions may also occur concurrently. For instance, when a person “racing a motorcycle”, he is also “riding and sitting on the motorcycle”. Inspired by that, we propose interaction correlation matching to infer other behaviors strongly correlated with the $top\text{-}1$ selected initial interaction action, as shown in Figure 4. When “race” is selected based on the highest similarity, we will also extract highly correlated actions, such as “ride”, “straddle”, “sit”, etc. Afterward, we amplify the similarity of the selected action and update the similarity vector. To be specific, for each row vector in $\bm{S}$, we employ a $top\text{-}1$ selection strategy to choose the initial interaction action with the highest image-text similarity. Furthermore, we amplify image-text similarity to highlight the similarity between these interaction actions. Detailed steps for applying interaction correlation matching to a single candidate image are:
@@
where $\bm{k}_{max}(\bm{S}_{i})$ represents the index of interaction action with the highest image-text similarity in the $i$th candidate image, the function $\mathbf{F}_{ICM}$ selects several other behaviors strongly correlated with $\bm{k}_{max}(\bm{S}_{i})$, $\bm{k}_{a}$ denotes a set of indexes associated with several interaction actions correlated with $\bm{k}_{max}(\bm{S}_{i})$, and $scale$ denotes the amplification factor.
@@
$\bm{mask}$ $=[m_{1},m_{2},\ldots,m_{j},\ldots,m_{N_{T}}]$
=$=\left\{\begin{matrix}0,j\notin\bm{k}_{o}\\ 1,j\in\bm{k}_{o}\end{matrix}\right.,j=1,2,…,N_{T},$ (10)
@@
where $\bm{o}_{i}$ represents the object category in the $i$th image. The function $\mathbf{F}_{PKM}$ selects the indexes of all interaction actions related to $\bm{o}_{i}$, denoted as $\bm{k}_{o}$. $m_{j}$ = {1, 0}: 1 indicates that the $j$th action is related to $\bm{o}_{i}$, not the other way around. Therefore, we retain interaction actions related to specific object categories after reducing interference from unlikely actions.
@@
$\bm{a}_{index}$
=$=\{j|(max\left(\bm{S}_{i}\right)-max\left(\bm{S}_{i}\right)\times\omega_{2})<\bm{S}_{ij}<max\left(\bm{S}_{i}\right),$
$\bm{S}_{ij}\in\bm{S}_{i}\},$ (13)
@@
IV Experiments
Datasets. The benchmark datasets, HICO-Det and V-COCO, are used to demonstrate the effectiveness of the proposed method. HICO-Det includes 47,776 images with 38,118 for training and 9,658 for testing, and covers 80 object categories, 117 action categories, and 600 distinct interaction types. V-COCO comprises 10,326 images with 5,400 for training and 4,964 for testing, featuring 80 object categories and 29 action categories, including 4 body actions without object interactions. Both datasets encompass a wide range of object and action categories, making them essential for evaluating HOI detection algorithms.
TABLE I: Performance comparisons on HICO-Det dataset. * represents the results given in  [26, 23].
Methods Backbone Source Default (mAPnormal-$\uparrow$) Know Object (mAPnormal-$\uparrow$)
Full Rare None-Rare Full Rare None-Rare
fully supervised (using $\langle$“human”, “interaction”, “object” $\rangle$ labels)
InteractNet [13] ResNet-50-FPN CVPR 2018 9.94 7.16 10.77 - - -
iCAN [45] ResNet-50 BMCV 2018 14.84 10.45 16.15 16.26 11.33 17.73
PMFNet [46] ResNet-50-FPN ICCV 2019 17.46 15.56 18.00 20.34 17.47 21.20
DJ-RN [47] ResNet-50 CVPR 2020 21.34 18.53 22.18 23.69 20.64 24.60
IDN [48] ResNet-50 NeurIPS 2020 23.36 22.47 23.63 26.43 25.01 26.85
HOTR [49] ResNet-50 CVPR 2021 25.10 17.34 27.42 - - -
QPIC [40] ResNet-101 CVPR 2021 29.90 23.92 31.69 32.38 26.06 34.27
HRNet [20] ResNet-152 TIP 2021 21.93 16.30 23.62 25.22 18.75 27.15
MSTR [50] ResNet-50 CVPR 2022 31.17 25.31 33.92 34.02 28.83 35.57
DisTr [51] ResNet-50 CVPR 2022 31.75 27.45 33.03 34.50 30.13 35.81
RCL [52] ResNet-50 CVPR 2023 32.87 28.67 34.12 35.52 30.88 36.45
GEN-VLKT [9] ResNet-50 CVPR 2022 33.75 29.25 35.10 36.78 32.75 37.99
PBLQG [11] ResNet-50 TIP 2023 31.64 26.23 33.25 34.61 30.16 35.93
SG2HOI [12] ResNet-50 TIP 2023 33.14 29.27 35.72 35.73 32.01 36.43
TED-Net [10] ResNet-50 TCSVT 2024 34.00 29.88 35.24 37.13 33.63 38.18
weakly+ supervised (using $\langle$“interaction”, “object” $\rangle$ labels)
Explanation-HOI* [53] ResNeXt101 ECCV 2020 10.63 8.71 11.20 - - -
MAX-HOI [24] ResNet101 WACV 2021 16.14 12.06 17.50 - - -
Align-Former [44] ResNet-101 arXiv 2021 20.85 18.23 21.64 - - -
PPR-FCN* [25] ResNet-50 ICCV 2017 17.55 15.69 18.41 - - -
Weakly-HOI [23] ResNet-50 CVPR 2023 19.26 - - - - -
PGBL [26] ResNet-50 ICLR 2023 22.89 22.41 23.03 - - -
Ours ResNet-50 - 24.57 21.45 25.51 26.52 23.64 27.38
weakly supervised (using $\langle$“interaction” $\rangle$ labels)
SCG* [54] ResNet-50 ICCV 2021 7.05 - - - - -
Weakly-HOI [23] ResNet-50 CVPR 2023 8.38 - - - - -
Ours (no labels) ResNet-50 - 16.96 16.26 17.17 18.89 18.11 19.12
TABLE II: Performance comparisons on V-COCO dataset. * represents the results given in  [23]. Weakly-HOI† is trained via video captured labels.

### v2

# FreeA: Plug-and-play Human-object Interaction Detection with Free Labeling

@@ Abstract
Recent human-object interaction (HOI) detection methods depend on extensively annotated image datasets, which require a significant amount of manpower. In this paper, we propose a novel self-adaptive, language-driven HOI detection method, termed FreeA. This method leverages the adaptability of the text-image model to generate latent HOI labels without requiring manual annotation. Specifically, FreeA aligns image features of human-object pairs with HOI text templates and employs a knowledge-based masking technique to decrease improbable interactions. Furthermore, FreeA implements a proposed method for matching interaction correlations to increase the probability of actions associated with a particular action, thereby improving the generated HOI labels. Experiments on two benchmark datasets showcase that FreeA achieves state-of-the-art performance among weakly supervised HOI competitors. Our proposal gets +13.29 (159$\uparrow$) mAP and +17.30 (98$\uparrow$) mAP than the newest “Weakly” supervised model, and +7.19 (28$\uparrow$) mAP and +14.69 (34$\uparrow$) mAP than the latest “Weakly+” supervised model, respectively, on HICO-DET and V-COCO datasets, more accurate in localizing and classifying the interactive actions. The source code will be made public.
1 Introduction
@@
Whether one-stage [39, 20, 34, 32] or two-stage [9, 21, 11, 31] HOI detection models, they predominantly rely on computationally heavy training and requires extensive-annotation datasets (Figure 1(a) and Figure 1(b)). Taking the train set of the HICO-Det dataset as an example, even for weakly or weakly+ supervised approaches, it still needs to annotate 117,871 interaction labels from (117,871$\times$600) or (117,871$\times$23) potential combinations. It is quite resource-intensive. Current weakly supervised HOI models can be divided into two folds (as shown in Figure 1(b)): weakly+ using $\langle$“interaction”, “object”$\rangle$ label  [13, 30], e.g., eat-banana, and weakly with only $\langle$“interaction”$\rangle$ labels, e.g., eat. Both still require abundant annotations for large-scale datasets to achieve satisfactory performance.
Another sub-direction of HOI detection is zero-shot HOI detection [7, 8, 20, 22], where models train on a subset of annotated data but test to unseen interaction categories. Although it aims to identify novel samples, it still depends on manually annotated training data.
Inspired by the effectiveness of the text-image matching model, e.g., CLIP [23], BLIP [14], and BLIP2 [15], in accurately pairing with a given image, FreeA is designed to accomplish HOI detection without relying on manual labels (Figure 1(c)). The FreeA mainly comprises three-folds: candidate image construction (CIC), human-object potential interaction mining (PIM), and human-object interaction inference (HII). In the phase of CIC, FreeA is plug-and-play, applying existing object detection methods for all potential instances localization, and utilizes spatial denoising and pairing techniques to establish candidate interaction pairs within the image. The PIM module extensively leverages the adaptability of the text-image model in the target domain to align the high-dimensional image features with HOI interaction templates, which generates the similarity vectors of candidate interaction relationships. Then, the HII module combines the resulting similarity vectors with a priori HOI action masks to mitigate interference from irrelevant relationships, and augments the likelihood of specific actions via the proposed interaction correlation matching method. Moreover, an adaptive threshold in HII generates HOI labels for training dynamically.
Our key contributions are summarized as three-folds:
@@
    We propose a novel HOI detection method, namely FreeA, that automatically generates HOI labels. To the best of our knowledge, it is the first framework to successfully achieve HOI detection tasks without the need for manual labeling.
@@
    HOI detection includes various interactions among multiple instances. Three key challenges are required to tackle when generating labels, namely, multiple actions selection, filtering out irrelevant actions, and refining the coarse text-image matching’s coarse labels. To address them, three corresponding modules are presented that significantly improve the effectiveness of the localization and classification of the interaction.
@@
    A broad variety of experiments are conducted to demonstrate the remarkable results of the proposal on the HOI detection task, and ours performs the best among all weakly+, weakly, and fully supervised HOI models.
2 Related Work
@@
Supervised HOI Detection. Supervised HOI models train their networks with the help of manually annotation $\langle$“human”, “object”, “action”$\rangle$. The two-stage methods first use pre-trained object detection network [24, 5] to detect humans and objects, and then pair them one by one into the interactive discrimination network to achieve HOI detection [9, 21, 11, 31]. However, it is pretty inefficient for one-to-one pairing between humans and objects [19]. To address that, one-stage HOI detection based on transformer is gradually developed via end-to-end solution  [39, 20, 32, 34]. In addition, it has been verified that text information enables to improve the HOI detection performance [18, 20]. Current HOI approaches, e.g., GEN-VLKT [20] and TED-Net [32], took use of CLIP [23] to train the encoding network via image-text pairs.
@@
Figure 2: Method overview. Starting from an existing HOI model, we apply the candidate image construction to extract humans and objects by detection and segmentation, and establish one-to-one human-object pairing. The human-object potential interaction mining module gets initial HOI interaction labels from candidate image pairs, and uses text-image matching model for domain adaptation. The human-object interaction inference module further refines these interaction labels by using $a$ $prior$ knowledge-based mask to eliminate implausible actions and using the interaction correlation matching method to enhance relevant action similarity. Finally, HOI labels are generated for model training through a dynamic threshold selector and interaction action filter.
@@
Weakly Supervised HOI Detection. Weakly supervised HOI detection generally uses image-level interaction labels for training [10, 13, 28]. They can be divided into twofolds: “weakly+” with $\langle$“interaction”, “object”$\rangle$ annotations [10, 13, 30], and “weakly” with only $\langle$“interaction”$\rangle$ annotations [28]. MX-HOI [13] proposed a momentum-independent learning framework using weakly+ supervised ($\langle$“interaction”, “object”$\rangle$). Align-Former [10] proposed an “align layer” to achieve pseudo alignment for training based on transformer architecture. Nevertheless, these methods suffer from noisy human-object association and ambiguous interaction types [30]. Therefore, Weakly HOI-CLIP [30] used CLIP as an interactive prompt network and HOI knowledge to enhance interaction judgment at the instance level. VLHOI [28] applied a language model to query for unnecessary interaction pairing reduction with image-level $\langle$“interaction”$\rangle$ annotations. Both of them, however, are still dependent on pre-annotated datasets at the cost of manpower. Unlike the latest methods [28] and [30], we propose a method for HOI detection that does not require manual annotation.
@@
Zero-shot HOI Detection. The goal of zero-shot HOI detection is to train on a subset of labels and test using another set of unseen labels to detect interactions that were not encountered during training. Many methods [7, 8, 25, 27, 20, 22, 3] are investigated to handle zero-shot HOI detection. Shen et al. [25] pioneered the application of zero-shot learning methods to address the long-tail problem in HOI detection. GEN-VLKT [20] is a simple yet effective framework that utilizes CLIP for knowledge transfer [22, 28], thereby discovering unknown samples. The introduction of zero-shot learning enhances the adaptability of these methods to real-world scenarios. However, these methods still require manually annotated complete HOI labels.
3 Method
@@
As shown in Figure 2, we propose a novel plug-and-play HOI framework, namely, FreeA, to reduce the requirements of annotations. The proposed framework includes candidate image construction, human-object potential relationship mining, and human-object relationship inference. Details are introduced below.
3.1 Candidate Image Construction
@@
To achieve automatic labels generation, we need to accurately localize humans and objects before, where Yolov8111https://github.com/ultralytics/ultralytics model is used for localization. Given an input image $\bm{I}$, we obtain a collection of instance bounding boxes $\bm{B}=\{(\bm{b}_{i}|i=1,2,...,N)\}$, where $\bm{b}_{i}=(x_{i},y_{i},W_{i},H_{i},c_{i})$ from Yolov8, with $x_{i}$ and $y_{i}$ representing the center coordinates of the $i$th bounding box. $W_{i}$ and $H_{i}$ are the width and height of the bounding box, and $c_{i}$ denotes the category of the instance within the box. Here, $N$ represents the number of detected instances.
@@
Subsequently, we trim the images using the bounding boxes and then pair humans with objects to create candidate images. $N_{I}$ candidate images are obtained, where $N_{I}=N_{h}\times N_{o}$, and $N_{h}+N_{o}=N$. $N_{h}$ and $N_{o}$ represent the number of humans and objects, respectively. To eliminate interference from redundant instances and background information, we apply instance segmentation to assist the PIM module in focusing on the interest of interactions.
3.2 Human-object Potential Interaction Mining
@@
It aims to mine potential relationships between humans and objects. We used the CLIP model to initialize the text encoder and image encoder. It is used to transfer knowledge from the source domain to the target domain. Then by computing cross-modal similarity, text-image pairs are matched.
@@
Image Encoder. The image encoder $\mathbf{F}_{IE}$ is employed to process the set of $N_{I}$ candidate images, with the result of image encoding $\bm{I}_{E}\in\mathbb{R}^{N_{I}\times C_{IE}}$. The $\bm{C}_{IE}$ denotes the dimension of the image encoder, and $\bm{I}_{B}$ represents the collection of candidate images. We have:
@@
Text Encoder. We start with a text template creation, denoted as $\bm{T}_{1}$, in the format “a photo of a person verb-ing an object”. For example, the triplet $\langle$ “human”, “ride”, “motorcycle”$\rangle$ is transformed into “a photo of a person riding a motorcycle”. After that, $\bm{T}_{1}$ is input into the text encoder $\mathbf{F}_{TE}$, leading to a text information matrix $\bm{T}_{E1}$ $\in\mathbb{R}^{N_{T}\times C_{TE}}$. $N_{T}$ denotes the number of textual queries, i.e., the number of HOI interaction categories. For the HICO-Det dataset, $N_{T}=600$; for V-COCO, $N_{T}=284$. $C_{TE}$ represents the dimensionality of the encoded textual features.
To emphasize the importance of verbs in HOI relationships, another type of text template $\bm{T}_{2}$ in the format “a photo of a person verb-ing”, has been constructed. $\bm{T}_{1}$ and $\bm{T}_{2}$ are distinct text templates for different HOI actions, corresponding to information matrices $\bm{T}_{E1}$ and $\bm{T}_{E2}$, respectively. They are formulated as:
@@
3.3 Human-object Interaction Inference
@@
Interaction Correlation Matching (ICM). If a certain action occurs, other actions may also occur concurrently. For instance, when a person “racing a motorcycle", he is also “riding and sitting on the motorcycle". Inspired by that, we propose interaction correlation matching to infer other behaviors strongly correlated with the $top\text{-}1$ selected initial interaction action, as shown in Figure 3. When “race" is selected based on the highest similarity, we will also extract highly correlated actions, such as “ride", “straddle", “sit", etc. The actions that are highly relevant to the target action
Figure 3: Details of the interaction correlation matching method.
are precomputed, and selecting the related actions is similar to a dictionary operation. Afterward, we amplify the similarity of the selected action and update the similarity vector. To be specific, for each row vector in $\bm{S}$, we employ a $top\text{-}1$ selection strategy to choose the initial interaction action with the highest image-text similarity. Furthermore, we amplify image-text similarity to highlight the similarity between these interaction actions. Detailed steps for applying interaction correlation matching to a single candidate image are:
@@
where $\bm{k}_{max}(\bm{S}_{i})$ represents the index of interaction action with the highest image-text similarity in the $i$th candidate image, the function $\mathbf{F}_{ICM}$ is a preprocessed dictionary that outputs a set of actions highly related to a given action category. $\bm{k}_{a}$ denotes a set of indexes associated with several interaction actions correlated with $\bm{k}_{max}(\bm{S}_{i})$, and $scale$ denotes the amplification factor.
@@
$\bm{mask}=[m_{1},m_{2},\ldots,m_{j},\ldots,m_{N_{T}}]=\left\{\begin{matrix}0,j\notin\bm{k}_{o}\\ 1,j\in\bm{k}_{o}\end{matrix}\right.,j=1,2,…,N_{T},$ (8)
@@
where $\bm{o}_{i}$ represents the object category in the $i$th image. The function $\mathbf{F}_{\text{PKM}}$ is a preprocessed dictionary that outputs all possible actions, denoted as $\bm{k}_{o}$, associated with a given object category. $m_{j}$ = {1, 0}: 1 indicates that the $j$th action is related to $\bm{o}_{i}$, not the other way around. Therefore, we retain interaction actions related to specific object categories after reducing interference from unlikely actions.
Table 1: Performance comparisons on HICO-Det dataset. * represents the results given in  [30, 28]. Swin-L represents the Swin-Transformer Large model.
[…]
Full Rare None-Rare
fully supervised (using $\langle$“human", “interaction", “object" $\rangle$ labels)
InteractNet [6] Faster R-CNN ResNet-50-FPN CVPR 9.94 7.16 10.77
iCAN [4] Faster R-CNN ResNet-50 BMCV 14.84 10.45 16.15
PMFNet [29] Faster R-CNN ResNet-50-FPN ICCV 17.46 15.56 18.00
DJ-RN [16] Faster R-CNN ResNet-50 CVPR 21.34 18.53 22.18
IDN [17] Faster R-CNN ResNet-50 NeurIPS 23.36 22.47 23.63
QPIC [26] DETR ResNet-101 CVPR 29.90 23.92 31.69
RCL [12] DETR ResNet-50 CVPR 32.87 28.67 34.12
TED-Net [32] DETR ResNet-50 TCSVT 34.00 29.88 35.24
PViC [36] DETR ResNet-50 ICCV 34.69 32.14 35.45
PViC [36] DETR Swin-L ICCV 44.32 44.61 44.24
RLIPv2 [34] DETR Swin-L ICCV 43.23 45.64 45.09
weakly+ supervised (using $\langle$“interaction", “object" $\rangle$ labels)
Explanation-HOI* [1] Faster R-CNN ResNeXt-101 ECCV 10.63 8.71 11.20
MAX-HOI [13] Faster R-CNN ResNet-101 WACV 16.14 12.06 17.50
Align-Former [10] DETR ResNet-101 arXiv 20.85 18.23 21.64
PPR-FCN* [37] R-FCN ResNet-50 ICCV 17.55 15.69 18.41
Weakly HOI-CLIP [30] Faster R-CNN ResNet-50 ICLR 22.89 22.41 23.03
Ours Faster R-CNN ResNet-50 - 24.14 20.91 24.09
OpenCat [38] Faster R-CNN ResNet-101 CPVR 25.82 24.35 26.19
Ours Faster R-CNN ResNet-101 - 27.02 24.49 28.37
Ours DETR ResNet-50 - 24.33 21.13 24.77
Ours Yolov8 ResNet-50 - 24.57 21.45 25.51
Ours Yolov8 Swin-L - 33.01 32.13 33.27
weakly supervised (using $\langle$“interaction" $\rangle$ labels)
SCG* [35] Faster R-CNN ResNet-50 ICCV 7.05 - -
VLHOI [28] Faster R-CNN ResNet-50 CVPR 8.38 - -
Ours (no use labels) Faster R-CNN ResNet-50 - 15.85 15.11 16.69
Ours (no use labels) Yolov8 ResNet-50 - 16.96 16.26 17.17
Ours (no use labels) Yolov8 Swin-L - 21.67 23.69 21.06
@@
$\bm{a}_{index}=\{j|(max\left(\bm{S}_{i}\right)-max\left(\bm{S}_{i}\right)\times\omega_{2})<\bm{S}_{ij}<max\left(\bm{S}_{i}\right),\bm{S}_{ij}\in\bm{S}_{i}\},$ (11)
@@
4 Experiments
Datasets. The benchmark datasets, HICO-Det and V-COCO, are used to demonstrate the effectiveness of the proposed method. HICO-Det includes 47,776 images and covers 80 object categories, 117 action categories, and 600 distinct interaction types. V-COCO comprises 10,346 images, featuring 80 object categories and 29 action categories, including 4 body actions without object interactions.
Table 2: Performance comparisons on V-COCO dataset. * represents the results given in  [28]. VLHOI† is trained via video-captured labels. The object detector of our is Yolov8.